# Hype-Adjusted Probability Measure for NLP Stock Return Forecasting

### v2

@@ 5.1 Model Results
Figure 8: Model Validation Accuracy Comparison for Market Return and Volatility Forecasting
@@
Figure 8 demonstrates the progression of accuracy across different models. For predicting market return direction, accuracy improved from $51.7\%$ in the baseline model to $70.0\%$ with a Na"ive model using adjusted scores, and further to $78.3\%$ with the best-optimized scores. Similarly, for volatility direction, accuracy increased from $53.8\%$ in the baseline model to $72.1\%$ and $75.0\%$ in the corresponding models. These substantial improvements validate the effectiveness of our framework.
@@
The observed accuracy increase of $+8.3\%$ for market return direction and $+2.9\%$ for volatility direction highlights the refined precision achieved through optimized sentiment modeling. This improvement translates to more reliable predictions of market trends, a critical factor for decision-making in high-volatility scenarios.
While there remains potential for refining the Naïve models, the current results validate the advantages of the proposed adjustments. Notably, even a Naïve scores model demonstrates promise as a foundational tool for forecasting prospects, showcasing the adaptability and robustness of the approach.
The primary contribution of this paper lies not in detailing machine learning model implementations but in the introduction of a novel probability measure that innovatively connects sentiment analysis with market dynamics.
Beyond validating improvements in financial forecasting, these results pave the way for broader theoretical applications. In the subsequent section 4, we introduce the hype-adjusted probability measure designed to quantify and integrate market hype, thereby extending the framework beyond traditional sentiment analysis.
5.2 Future Work
Building upon the introduction of the hype-adjusted probability measure $\mathbb{P}^{a}$, our future work aims to extend this concept into a more comprehensive probability theory framework. By leveraging insights from the validation and calibration methods of the Mathematical Model lacking close-form solutions in recent work [cao2024theoreticalempiricalvalidationheston], we plan to develop an algorithm that computes the Radon-Nikodym derivative or Girsanov theorem set function and optimize constant and time-dependent parameters based on the forecast output of various models.
This approach will refine the sentiment model by systematically addressing data biases, such as political leanings or disproportionate media attention given to certain stocks (e.g., Nvidia receiving excessive coverage compared to its sector weight). The algorithm will also incorporate parameters like variance and precision in the loss function to improve prediction accuracy, assuming identical training models and data dimensions.
Future work will also explore integrating sentiment-guided adversarial learning frameworks, such as those presented in [sentiment-guided2023], to enhance the ability of the model to adapt to dynamic market conditions. By leveraging these advanced techniques, we aim to better capture the nuanced interplay between market hype and fundamental indicators, improving the generalizability of the hype-adjusted probability measure.
By aligning sentiment scores with security prices as time approaches maturity, this method will create a more objective measure to correct biases from the data sources mentioned earlier. The goal is to ensure that the sentiment-driven predictions are more reflective of market fundamentals, especially when sentiment diverges significantly from the stock’s actual influence in the sector.
Additionally, future research will explore the incorporation of further indicator functions into the model. For instance, introducing non-zero thresholds, $\mathbbm{1}_{|Sent_{d}-Sent_{d-1}|>c}$, could enhance the model by accounting for market behavior where significant impacts occur only when the sentiment change exceeds a certain threshold, $c$. This would provide a more nuanced understanding of how sharp sentiment shifts influence volatility and pricing, reinforcing the utility of the hype-adjusted probability measure $\mathbb{P}^{a}$ in forecasting financial markets.
Further analysis of bias or subjectivity could also be conducted using tools like the Eikon Data API, as discussed earlier in the paper. This will help deepen our understanding of the interplay between sentiment, volatility, and pricing, and further validate the efficacy of the hype-adjusted probability measure in real-world applications.
Our next paper is aiming to derive and formalize the adjusted volatility and option pricing proposals in section 4.2.3
@@ 6 Conclusion
In this paper, we presented an improved NLP approach for forecasting stock return and volatility based on the introduction of the new sentiment equation, the “3 weights", and the hype-adjusted probability measure, $\mathbb{P}^{a}$.
Our findings provide quantitative support for the market adage "buy on rumors, sell on news," as we confirm a stronger correlation between negative news sentiment and market volatility, as well as price declines. This highlights the asymmetrical influence of negative sentiment on the market compared to positive sentiment.
Our model evaluates text tokens with greater precision by integrating historical data, effectively capturing the influence of correlations and weights across subsets. A key advantage is its low sensitivity to the precision of the initial token scoring method, making it robust across varied data sources. Figure 8 demonstrates the enhanced performance of our sentiment model, showing higher accuracy in predicting market responses compared to baseline models.
Most importantly, we introduced the concept of a new probability measure, the hype-adjusted probability measure, $\mathbb{P}^{a}$, and outlined potential avenues for future research, including deriving new option pricing formulas, refining sentiment weighting parameters, and expanding applications within and beyond market forecasting. [^1] We are over the limit of our knowledge.
@@ 7 Acknowledgements
Lastly, we want to express our heartfelt gratitude for the insightful comments and constructive suggestions of the reviewers.
@@ Appendix A Appendix A: Ticker News Weight Table
For the table below, Close Price refers to the price per share at market close in U.S. dollars; Capital is reported in millions of dollars; Capital Weight % represents the company’s market capitalization as a percentage of the entire sector; and News Weight % indicates the proportion of news coverage the company receives relative to the whole sector.

### v4

@@ 5.1 Results
Figure 8: Model Accuracy Comparison for Return and Volatility Forecasting (discussed below)
We choose volatility and market return directions (increase or decrease) as the target of prediction, which aligns with the observation results from sections 3.4 and 4.1.
@@
Figure 8 demonstrates the progression of accuracy across different models. In the prediction of market return direction, validation accuracy improves from $51.7\%$ in the baseline model to $70.0\%$ with a LDA model (as presented by Deveikyte et al [deveikyte2022]) using adjusted scores, and further to $78.3\%$ with the optimized scores. Similarly, for volatility direction, accuracy increases from $53.8\%$ in the baseline model to $72.1\%$ and $75.0\%$ in the corresponding models. These substantial improvements validate the effectiveness of our methodology.
5.2 Implications
@@
The observed accuracy increase of $+8.3\%$ for market return direction and $+2.9\%$ for volatility direction highlights the refined precision achieved through optimized sentiment modeling. This improvement translates into more reliable predictions of market trends, a critical factor for decision-making during high-volatility scenarios and market crises.
The results have demonstrated significant enhancement of the NLP forecasting approach with the adjusted scores and thus further support the validation of a hype-adjusted probability measure.
This research extends classical finance theory by integrating sentiment adjustments into probabilistic modeling, bridging asset pricing and NLP methodologies. It formalizes sentiment as a quantitative factor, enriching models of asset returns and volatilities while addressing bias, memory, and directional shifts in market dynamics. The work represents a foundational interdisciplinary contribution, connecting finance, probability theory, and machine learning.
The model enhances forecasting accuracy for investment strategies by correcting sentiment biases, enabling better-informed decision-making. Policymakers can leverage these insights to understand sentiment-driven market behavior and regulate excessive hype or misinformation. Additionally, the framework aids in risk management and thematic investing by identifying overhyped sectors and improving resource allocation.
5.3 Limitations and Future Research
Note that we do not claim the uniqueness of the hype-adjusted probability measure $\mathbb{P}^{a}$ we are proposing, for two sets of reasons:
1. The way we define “hype" can vary based on the approach used to link news to market sentiment. For instance, other researchers may define hype using different NLP techniques, sentiment scoring models, or thresholds. There is no canonical way of defining hype, and our chosen approach is just one of many possibilities to bridge news sentiment with market behavior.
2. Moreover, there could be multiple valid constructions of $\mathbb{P}^{a}$ that align with the sentiment adjustments.
The flexibility in both defining hype and constructing the adjusted measure ensures the generality of the approach, allowing it to be adapted to various datasets and market contexts.
Regarding our potential future research, one avenue could be to derive and formalize a hype-adjusted volatility and option pricing under a hype-adjusted probability measure. Future work can explore integrating sentiment-guided adversarial learning frameworks, such as Long Short-Term Memory (LSTM) networks and generative adversarial networks (GANs) presented, for instance, by Zhang et al in [sentiment-guided2023], to enhance the ability of the model to adapt to dynamic market conditions. Further analysis of bias or subjectivity can also be conducted using tools like the Eikon Data API, as discussed earlier in the paper.
An important point to keep in mind is that market participants do not only trade on news. Some trade on the basis of technical analysis, such as moving averages (for commodities but for equities as well); a small category of market participants, called arbitragers, trade only if they have identified a strict arbitrage opportunity; some major hedge funds try to recognize “statistical arbitrage" patterns. Lastly, a category of fundamental analysts uses Capital Structure and Earnings as trading signals. All these strategies interact with each other, leading to market prices whose exact formation is beyond the scope of this paper.
@@ 6 Conclusion
In this paper, we present an improved NLP approach for forecasting stock return and volatility based on a hype-adjusted probability measure, $\mathbb{P}^{a}$.
Besides validating improvements in financial forecasting, these results lay the foundation for broader theoretical applications. A hype-adjusted probability measure is introduced to quantify and integrate market hypes, extending the framework beyond classical sentiment analysis.
Our model evaluates text tokens with greater accuracy, effectively capturing the influence of correlations and weights across subsets. A key advantage is its low sensitivity to the accuracy of the initial token scoring method, making it robust across varied data sources. The enhanced performance of our sentiment model is exhibited in the key result figure which shows the higher accuracy obtained in predicting market responses compared to baseline models.
Lastly, our hype-adjusted probability measure can arguably be valued as a theoretical bridge between the probabilistic finance of Asset Pricing and NLP prediction in Finance, two fields which are likely to intersect more and more in the near future.
@@ Appendix A Appendix A: Ticker News Weight Table
For the table below, Close Price refers to the price per share at market close in U.S. dollars; Capital is reported in millions of dollars; Capital Weight % represents the company’s market capitalization as a percentage of the entire sector’s capitalization; and News Weight % indicates the proportion of news coverage the company receives relative to the whole sector.
Note, in the first row, the remarkable difference between columns 3 and 4.